OpenStudy (nwolf):

Aisha just received a $50 paycheck from her new job. She spends some of it buying music online and saves the rest in a bank account. Her savings is given by f(x) = 50-1.30x, where x is the number of songs she downloads.

OpenStudy (anonymous):

whats the question?

OpenStudy (nwolf):

f(x)=50-1.30x

OpenStudy (anonymous):

what about it, its an equation are you asking i its the correct equation for the given information?

OpenStudy (nwolf):

Find F(18) but i am also not sure what they are asking

OpenStudy (anonymous):

they are asking find (50-1.30x) when x=18 f(x) gives you 50-1.30x f(18) gives you50-1.30(18)

OpenStudy (anonymous):

Yep, the wordy's weird!

OpenStudy (nwolf):

Ok now it is asking what does f(18) represent? Would that be the number of songs downloaded. If I did the problem correctly from 50-1.30(18) I got 26.6 does that sound right?

OpenStudy (anonymous):

18 does represent the number of songs she purchased, do you understand what the 26.6 represents?

OpenStudy (nwolf):

no is that the price per song

OpenStudy (anonymous):

@nwolf focus on the question?

OpenStudy (anonymous):

"Her savings is given by f(x) = 50-1.30x, ..."

OpenStudy (anonymous):

You're working on the f(x), so what is its purpose?

OpenStudy (anonymous):

@nwolf Do you follow?

OpenStudy (nwolf):

no the F(x) messes me up.

OpenStudy (anonymous):

if x=number of songs purchased what would you be figuring out using that equation? f(x) is just notation its just a number dont let it scare you

OpenStudy (nwolf):

so can you tell me what the 26.6 stands for or does that matter?

OpenStudy (anonymous):

it represents the amount that she saved

OpenStudy (anonymous):

That's why the sentence starts with "Her saving ..."

OpenStudy (nwolf):

oh

OpenStudy (anonymous):

f(x) is function to find her saving :)

OpenStudy (nwolf):

ok just to see if I understand this right, she saves 26.6 but if she wants to buy songs and still save 40$ and she buys 18 songs what is the price per song? Do I take 40 and divide by 18 and that should give me the price per song??

OpenStudy (anonymous):

The price per song is $1.30!

OpenStudy (anonymous):

It's a default value, the variable is the number of songs!

OpenStudy (nwolf):

oh my!! not I lost myself all over again.. so at 1.30 she can buy 7 songs since she wants to save 40$$ is that right?

OpenStudy (anonymous):

Yep, look like you set it straight now :)

OpenStudy (anonymous):

If you want us to help you easily, you need to post entirely problem, otherwise we have to guesssing around, then mislead you. That's never our intend!

OpenStudy (anonymous):

Is it crystal clear about the purpose of f(x) and what x stands for now?

OpenStudy (nwolf):

then can you tell me what is mant by what does f(18)) represent?

OpenStudy (anonymous):

f(x) = f(18) means x = 18 songs

OpenStudy (nwolf):

this part I am kinda lost on.

OpenStudy (anonymous):

f(18) = 50 - 1.3 * 18 1.3 * 18 is the cost of 18 songs => 50 - (1.3 * 18) is her saving left over after the cost of 18 songs which f(18)

OpenStudy (nwolf):

what I am lost on is if 18 the the amount of songs then she can't buy 18 because she only wants to spend 10$

OpenStudy (anonymous):

Since you don't post the whole problem, we can't understand where you lost?

OpenStudy (anonymous):

On your post, there's no number 18, 10, 40,.. so we don't know where do them come from?

OpenStudy (nwolf):

ok let me try again to post it and see Aisha just received a $50 paycheck from her new job. She spends some of it buying music online and saves the rest in a bank account. Her savings is given by f(x) = 50-1.30x, where x is the number of songs she downloads. a. Find f(18). What does this value (the answer to f(18)) represent? b. How many songs can Aisha buy if she wants to save $40?

OpenStudy (anonymous):

See, that's where we have trouble to help you!

OpenStudy (anonymous):

F(18) is the saving left over after purchase 18 songs!

OpenStudy (anonymous):

make f(x)=40 so 40=50-1.30x and solve for x

OpenStudy (anonymous):

@Mcronin, thank god you come back, I just don't know head and tail :(

OpenStudy (anonymous):

Because f(x) is to calculation for saving, so you set it equal to $40 That's how algebra works, it's sort of backward !

OpenStudy (anonymous):

@nwolf questions?

OpenStudy (nwolf):

so I ended up getting 7.6 does that sound correct?

OpenStudy (anonymous):

Yep, almost 8 songs!

OpenStudy (anonymous):

I means 7 songs!

OpenStudy (anonymous):

If you can understand throughly on problem, next time you can solve million problems :)

OpenStudy (nwolf):

cool thanks so much so 7 songs @ 1.30???

OpenStudy (anonymous):

Yep!

OpenStudy (nwolf):

wow thank god for math people!!!

OpenStudy (anonymous):

Any more problem?

OpenStudy (nwolf):

I do have another math (word problem that I am lost on)

OpenStudy (anonymous):

SEE, the keywork "Her saving .. f(x)" You totally missed it, because you're so intimidate by the letter f(x) :P

OpenStudy (anonymous):

Just post it up, I'm restart my PC in 5 min!

OpenStudy (nwolf):

Examine the rise of the price of milk from 1997 to 2008. The average price of a gallon of whole milk in January 1997 was $2.67, and in January 2008, the average price of a gallon of whole milk was $3.87 according to Consumer Price Index. Use the coordinates (1997, 2.67) and (2008, 3.87) to find the slope, or rate of change, between the two points. Describe and show how you arrived at your answer. Express your answer as a decimal rounded to the nearest hundredth. What does this answer mean the real world?

OpenStudy (anonymous):

(1997, 2.67), (2008, 3.87) Slope, rate of change: = (3.87 - 2.67)/(2008 - 1997) = $.109/ year

OpenStudy (anonymous):

The cost of 1 gallon milk keep increasing averagely $.11 each year from 1997 to 2008! It means the cost of living raising up yearly!